• smooth operation in general
  • problem with migration to pilot2+container 
    • farm drained after the migration. Pilots failed and exited before picking up real jobs, due to a mismatch between pilot2 and wrapper version, because harvester uses a special template for BNL. The templte issue needs to be taken care of first. Backed off to pilot1 for now.  
  • several racks of T1 computing farm is draining now, for scheduled electrical work in the data center late this week and early next week.
  • continue work on integration of GPU@BNL IC with PanDA